Solve each problem.
Layoffs looming. American Products plans to lay off 10% of its employees in its aerospace division and 15% of its employees in its agricultural division. If altogether 12% of the 3000 employees in these two divisions will be laid off, then how many employees are in each division?

Answer by checkley71(8403)

.1X+.15(3000-X)=.12*3000 OR .1X+450-.15X=360 OR -.05X=-90 OR .05X=90 OR X=90/.05 OR X=1800 OR AEROSPACE WORKERS & 3000-1800=1200 AGRICULTURAL WORKERS.